Centroid of Plane Area


Locate the centroid of the region ๐‘… bounded by the given curves. Sketch the graph.


1.) ๐‘…:๐‘ฆ=๐‘ฅ^2, ๐‘ฆ=2๐‘ฅ^2 โˆ’3๐‘ฅ

2.) ๐‘…:๐‘ฆ^2=4๐‘ฅ, ๐‘ฆ=4, ๐‘ฆโˆ’๐‘Ž๐‘ฅ๐‘–๐‘ 

3.) ๐‘…:๐‘ฅ=2๐‘ฆโˆ’๐‘ฆ^2, ๐‘ฆโˆ’๐‘Ž๐‘ฅ๐‘–๐‘ 

4.) ๐‘…:๐‘ฆ=2๐‘ฅ+1, ๐‘ฅ+๐‘ฆ=7, ๐‘ฅ=8


1
Expert's answer
2021-05-23T08:16:34-0400

1)"x=\frac{\iint\limits_D xdxdy}{\iint\limits_D dxdy}=\frac{\int\limits_0^3\int\limits_{2x^2-3x}^{x^2}xdydx}{\int\limits_0^3\int\limits_{2x^2-3x}^{x^2}dydx}=\frac{\int\limits_0^3(x^2-2x^2+3x)xdx}{\int\limits_0^3(x^2-2x^2+3x)dx}=\frac{6.75}{4.5}=1.5\\
y=\frac{\iint\limits_D ydxdy}{\iint\limits_D dxdy}=\frac{0.5*\int_0^3 ((x^2)^2-(2x^2-3x)^2)dx}{4.5}=\frac{0.5*16.2}{4.5}=1.8"



2

"x=\frac{\iint\limits_D xdxdy}{\iint\limits_D dxdy}=\frac{\int\limits_0^44x\sqrt{x}dx}{\int\limits_0^44\sqrt{x}dx}=\frac{256/5}{64/3}=2.4"

due to symmetry about the x-axis, y = 0


3

"x=\frac{\iint\limits_D xdxdy}{\iint\limits_D dxdy}=\frac{\int_0^1 2*x*\sqrt{1-x}dx}{\int_0^1 2*\sqrt{1-x}dx}=\frac{8/15}{4/3}=0.4"

due to symmetry with respect to y = 1, y = 1


4

this is a triangle

"x=(1/3)*(2+8+8)=6\\
y=(1/3)*(5 -1+17)=7"